OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 3.0 centimeters/minute while the area of the triangle is increasing at a rate of 2.0 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 11.5 centimeters and the area is 91.0 square centimeters?

OpenStudy (anonymous):

area of a triangle = 1/2 base* height

OpenStudy (anonymous):

differentiate this equation wrt time

OpenStudy (anonymous):

u can write A= 1/2 b* h

OpenStudy (anonymous):

can u differentiate this equation wrt time?

OpenStudy (anonymous):

what will the equation that i have to differentiate be? what numbers would i sub in?

OpenStudy (anonymous):

differentiate this equation A= 1/2 b * h A= area o the triangle b= base h= height

OpenStudy (anonymous):

supposes u have been given x= 2t^2 and i ask u to differentiate this wrt t(time)

OpenStudy (anonymous):

x is the distance traversed..and if u differentiate it then u'll get the speed...right?

OpenStudy (anonymous):

forget the substitution part for now...that will come later

OpenStudy (anonymous):

x= 2t^2 differentiate x wrt time

OpenStudy (anonymous):

x=4t?

OpenStudy (anonymous):

it'll be dx/dt= 4t

OpenStudy (anonymous):

differentiation of x wrt t= dx/dt

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

what's ur confusion?

OpenStudy (anonymous):

I gave u an equation x=t^2 and asked u to differentiate wrt t

OpenStudy (anonymous):

i got the equation for my problem to equal 2=(1/2)dB/dT * 11.5+dH/dT*(1/2)b

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

A= 1/2 b * h dA/ dt = 1/2( db/dt *h + b * dh/dt) 2= 1/2 ( db/dt *h + b*3)

OpenStudy (anonymous):

u have been given dh/dt= 3cm/minute

OpenStudy (anonymous):

now u have to find db/dt when h=11.5 and A=91 cm^2

OpenStudy (anonymous):

see this equation 2= 1/2 ( db/dt *h + b*3) for finding db/dt u need to know b and h..but u know h only so find b

OpenStudy (anonymous):

for finding b,use A= 91 cm^2

OpenStudy (anonymous):

b=.063107 ?

OpenStudy (anonymous):

A=91, h=11.5 then b=?

OpenStudy (anonymous):

b=15.82 cm

OpenStudy (anonymous):

2=1/2(db/dt) *11.5+15.82*3 ?

OpenStudy (anonymous):

2=1/2(db/dt) *11.5+1/2 * 15.82*3

OpenStudy (anonymous):

u left 1/2 in the 2nd term of RHS

OpenStudy (anonymous):

so find db/dt that will be the answer.

OpenStudy (anonymous):

i got -.778223 but its telling me its the wrong answer

OpenStudy (anonymous):

how did u come to know that it's wrong answer?

OpenStudy (anonymous):

due to -ve sign?

OpenStudy (anonymous):

the retriceinment is online to i entered the answer, and its telling me its wrong

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

assignment*

OpenStudy (anonymous):

enter this value db/dt=-3.77

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

no its coming up incorrect